### v1

@@ 3 The Zero Rule
In this section, we propose the Zero rule which works on the dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ without the objective function:
@@
3) Pivot $(r,s)$ and update the dictionary.
The Zero rule represents an advance beyond the Criss-Cross rule, no longer focusing on the objective function but on the entries of the coefficient matrix, which is the arrangement itself. Now, we prove the desirable properties of the proposed Zero rule. First, we show that given a dictionary, the Zero rule will only select a unique entry to pivot, which is obtained by direct calculus.
@@
Proof: Direct calculus.
@@
Next, we give a condition to test if an entry of a dictionary passes the valid reverse pivoting test. The criterion is that if $(s,r)$ passes the test, then applying the Zero rule in the new dictionary will automatically pinpoint $(r,s)$. Then, pivoting $(r,s)$ will go back to the original dictionary. This property is called reversibility, which is important to the vertex enumeration algorithm in terms of ensuring that no storage is needed, and no vertices are missed. Figure 3 presents an example to illustrate the Zero rule and its valid reverse, where blocks of different colors correspond to different sub-conditions taking effect.
@@
Lemma 3.1 (Index increasing).
Starting with any dictionary arrangement, when the Zero rule is repeated, the sequence of indices leaving the co-basis in each step increases.
Proof: Assume in the $l$-th step, the Zero rule is fixed at $(r_{l},s_{l})$ and yields the $l$-th dictionary $\mathbf{x}_{B_{l}}=\bar{A}\mathbf{x}_{N_{l}}$, where the Zero rule pinpoints $(r_{l+1},s_{l+1})$.
@@
Based on Lemma 3.1, one can prove that the number of pivot steps of the Zero rule is no more than $d$.
Proposition 3 (Termination in $d$ steps).
For any dictionary in an arrangement of $\mathbb{R}^{d}$, consecutively applying the Zero rule or its reverse will attain the terminal in at most $d$ steps.
Proof: According to Lemma 3.1, it can be concluded that each time the Zero rule is applied, the column it selects will be at least one position later compared to the previous step. In any dictionary in $\mathbb{R}^{d}$, there are a total of $d$ columns, excluding the column corresponding to $x_{g}$. This implies that the Zero rule will be repeated for at most of $d$ times.
@@
The property given in Proposition 3 is very important because most pivot rules, such as Bland’s rule and the Criss-Cross rule, converge in exponential steps. The exponential convergence of Bland’s rule is proven in [27, 34], while the Criss-Cross rule is shown by an example that requires at least $12^{n}-$ 1 steps for convergence [29]. Moreover, a better method, such as the Random-Facet pivot rule mentioned in [14, 20], takes at most a subexponential number of pivot steps. Therefore, the convergence speed of our Zero rule is very fast.
After studying the number of convergence steps used in the proposed Zero rule, we focus on the number of terminations among the arrangement. Will the pivot rule lead to multiple terminations? To address this question, let us first identify what “0” means in each dictionary.
Lemma 3.2.
@@
for all $i\in B$ and $j\in N,j\neq g$. Substituting Eq. (3) into $x_{i}=\sum\limits_{j\in N}\bar{a}_{ij}x_{j}$, we can derive the relationship between $\mathbf{c}_{i}$ and $\mathbf{c}_{j}$ as follows:
@@
From the above lemma, it can be seen that the proposed Zero rule is not an indexing game. Instead, it is deeply related to the properties of the arrangement itself, as represented by the dictionary. The relationship between zero and non-zero entries is the reason why we call our proposed rule the Zero rule. Now, we show that under the Zero rule, there only exists one termination.
Proposition 4 (Unique convergence).
Repeating the Zero rule from any dictionary in the arrangement will result in a unique termination.
Proof: Let $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ be the dictionary such that $N$ has attained the smallest lexicographical order among all co-basis in the system; note that the dictionary is a terminal of the Zero rule. Furthermore, suppose there exists another terminal of the Zero rule dictionary $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ and $\tilde{N}\neq N$, that is,
@@
Since $N$ is a co-basis in the arrangement, then $\{\mathbf{c}_{l}~{}|~{}l\in N_{0}\}\cup\{\mathbf{c}_{n_{j}}\}\subseteq\{\mathbf{c}_{l}\ ~{}|~{}\ l\in N\}$ is linearly independent, which contradicts with Eq. (7). Hence, in the first place, any dictionary with a co-basis not attaining the lexicographical minimum order is not the end of pivoting by the Zero rule. It implies that the termination of repeating this rule from all dictionaries among an arrangement is unique, which is the co-basis with the lexicographical minimum order in the arrangement.
@@
Connection to an open question. Propositions 3 and 4 might be related to an open question: “Is it possible to find a polynomial pivot rule for linear programming, i.e., a pivot rule which bounds the number of pivot steps by a polynomial function of the number of variables and the number of inequalities, or to prove that such a pivot rule does not exist”. This question is important [34], motivated by analyzing the complexity of the simplex method in linear programming. The diameter of a polytope provides a lower bound on the number of steps needed by the simplex method, although it is actually weaker than the so-called Hirsch conjecture in polyhedral theory. The Hirsch conjecture, formulated by W. M. Hirsch in 1957 in [22] and reported in the 1963 book of Dantzig [13], states that the edge-vertex graph of an $n$-facet polytope in $d$-dimensional Euclidean space has a diameter no more than $n-d$. This means that any two vertices of a polytope must be connected by a path of length at most $n-d$. However, this conjecture has been refuted by a counterexample presented in [30] which introduces a 43-dimensional polytope of 86 facets with a diameter exceeding 43. Additionally, a formal disproof is also outlined in [2].
However, this counterexample has no direct consequences for the analysis of the simplex method, as they do not rule out the possibility of a larger but still polynomial number of steps. While the proposed Zero rule is a polynomial pivot rule and is bounded by a linear function, it is on an arrangement instead of a polytope. Nevertheless, the Zero rule can shed some light on the open question. The Zero rule is based on the dictionary itself and goes without a ratio test. Thus, it does not inherently traverse between adjacent vertices or reflect the number of edges between vertices. If a ratio test could be incorporated or primal feasibility could be maintained within the Zero rule, it might serve as a polynomial pivoting rule on polytopes.
@@ 4 VE Algorithm Using the Zero Rule
In the above section, we introduce the Zero rule and its favorable properties. Here, we design the vertex enumeration algorithm with the Zero rule as its core technique. The designed algorithm inherits the beautiful framework in Avis and Fukuda [4] including the lex-min test and applying a rule and its valid reverse, among others. The core is we replace the Criss-Cross rule therein with the proposed Zero rule. Below we describe the proposed algorithm, elucidate its advantages, and conduct a detailed complexity analysis.
4.1 Algorithm Description
The pivoting method involves two main steps: initialization and search [4]. The initialization finds a vertex from a given arrangement and constructs an initial dictionary based on this vertex. Specifically, we take $d$ hyperplanes that intersect at a vertex. Then, we arrange the associated slack variables as co-basis variables and index them from $1$ to $d$. The rest slack variables will be basis variables and indexed from $d+1$ to $n$. According to Proposition 4, the so-initialized dictionary is the termination dictionary of all other dictionaries, as it is the lexicographically minimum.
Second, searching all dictionaries requires testing valid reverse for every entry of a dictionary until the if-and-only-if condition is satisfied. The search is based on the order of increasing index. When an entry fulfills the if-and-only-if condition, we pivot it and update the dictionary. Recursively conducting the reverse until the pivot rule is not fixed at any entry, which means that the algorithm arrives at the termination dictionary. At this time, as Figure 5 shows, we apply the Zero rule to go back to the previous dictionary and search the entry again. This will not cause repetition because the if-and-only-if condition can inform us of the previously visited entry. We can search for entries with higher indexes starting from the previously visited one.
Moreover, in the searching step, for each obtained dictionary, the lex-min test is done to verify if the dictionary has a lexicographically minimum basis. Only the basis of the dictionary that passes the test will output the vertex. This operation is to avoid producing the same vertex more than once.
By applying the Zero rule and its valid reverse, the searching path is a single spanning tree, and any two branches do not intersect, as Figure 5 shows. Each node in the spanning tree is guaranteed to be unique. However, Avis and Fukuda [4] did not provide the if-and-only-if condition, which may cause loops in the searching path. Moreover, by Proposition 4, there is only one dictionary that can not be applied to the Zero rule, and we do not need to store any intermediate dictionary. A detailed algorithm is presented as Algorithm $1$ in Figure 6.
Figure 5: Explination of the VE algorithm.
Since the Zero rule can converge in $d$ steps, one can set the initial dictionary to the 00-th layer. Each reverse pivot increments the layer by one, while each application of the Zero rule decrements the layer by one. When studying the dictionary of the $(d-1)$-th layer, we only need to determine whether the corresponding dictionary of the $d$-th layer satisfies the lex-min condition instead of examining the valid reverse. Then one can track back to the $(d-1)$-th layer. This can further save time. Therefore, our algorithm needs to save the layer number which is of $O(1)$ complexity.
Figure 6: The VE algorithm of an arrangement using Zero rule.
4.2 Advantage Analysis
@@
Since our focus is VE of an arrangement, we only compare with the Criss-Cross rule and Jensen’s general relaxed recursion that are applicable to an arrangement. The advantages of the Zero rule are as follows:
@@
Jensen’s general relaxed recursion requires partitioning a dictionary and analyzing it case by case. Its termination dictionary could be primal feasible, dual feasible, or optimal [21, 34], which makes it tedious for VE. As for the Criss-Cross rule, although it also needs to discuss two cases, it has only one type of termination dictionary, namely the optimal dictionary. This makes it better than Jensen’s general relaxed recursion, and it is also exactly the pivot rule used by Avis and Fukuda [4].
However, since the condition by Avis and Fukuda is not if-and-only-if, which means that after performing reverse pivoting, applying the Criss-Cross rule may not return to the original dictionary directly. As Figure 7 shows, when the reverse-applying operation about the Criss-Cross rule goes to a new entry, if this new entry has a smaller index, the algorithm goes into a loop. If this new entry has a larger index, the algorithm catastrophically skips some vertices. Both cases make the vertex enumeration fail. The advantage of using the if-and-only-if condition when testing valid reverse pivoting in vertex enumeration is that it guarantees one can return to the original dictionary without requiring additional storage. Since it is automatically known that the $(s,r)$ entry in the original dictionary has been pivoted, one can search for entries greater than $s$ and $r$ in further study.
Figure 7: The Criss-Cross rule may undergo loops and catastrophic skipping.
@@
Moreover, compared to the Zero rule, the Criss-Cross rule is not ideal in both initialization and valid reverse. During the initialization process, the Zero rule is more convenient in generating all dictionaries where the pivot rule cannot be applied. By Proposition 4, through our way of initialization, the resultant dictionary with co-basis $\{g,1,2,\cdots,d\}$ is always the unique terminal dictionary. On the other hand, using Criss-Cross rules requires setting an objective function and generating all optimal dictionaries, which adds an extra step to the initialization process.
When it comes to the valid reverse, the advantages of the Zero rule are more obvious. The if-and-only-if condition between the Zero rule and its valid reverse is established so that all vertices are surely enumerated. But the Criss-Cross rule has a risk of being failed such as going into loops or skipping vertices. On the other hand, according to Proposition 3, the number of pivot steps of the Zero rule is bounded by $d$, whereas the pivot step of the Criss-Cross rule is exponential [29, 34]. Using the Zero rule, all dictionaries can be arrived in at most $d$-step reversals. At the same time, since the Zero rule can reach the unique termination in $d$ steps, using the Zero rule can skip testing valid reverse on dictionaries at the $d$-th layer, whereas the Criss-Cross rule requires studying every dictionary.

### v2

@@ 3 The Zero rule
In this section, we propose the Zero rule which works on the dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ without the objective function. Let $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ be a dictionary without the objective function. Then the Zero rule is given in the following:
@@
Remark 1: The Zero rule refers to a selection process that does not impose any requirements for primal feasibility or dual feasibility. Consequently, the dictionary where the Zero rule can not select a proper position could be neither primal feasible nor dual feasible dictionary. For convenience, we refer to it as a terminal dictionary, or simply as the terminal. The existence of the terminal dictionary is trivial. Note that any Zero pivot removes $s$ from the co-basis and introduces $r$ into it. Since $r<s$, after the Zero pivot, the new co-basis has a smaller lexicographical order. Therefore, the dictionary equipped with the smallest lexicographical order co-basis automatically becomes a terminal dictionary.
Now, we prove the desirable properties of the proposed Zero rule. First, we show that given a dictionary, the Zero rule can only select a unique pair of variables to pivot, which is obtained by direct computation.
@@
Proof: Direct calculation.
@@
Next, in order to have a better understanding of Zero pivots, we give an if-and-only-if condition to test if a pivot on a dictionary is a valid reverse Zero pivot. Figure 2 presents an example to illustrate the Zero rule and its valid reverse, where blocks of different colors correspond to different sub-conditions taking effect.
@@
Lemma 1 (Index increasing).
Starting with any dictionary in any arrangement, when the Zero pivot is repeated, the sequence of indices leaving the co-basis in each step increases.
Proof: Assume in the $l$-th step, the Zero rule select $(r_{l},s_{l})$ and yields the $l$-th dictionary $\mathbf{x}_{B_{l}}=\bar{A}\mathbf{x}_{N_{l}}$, where the Zero rule pinpoints $(r_{l+1},s_{l+1})$.
@@
Based on Lemma 1, one can prove that the number of Zero pivots recursively is no more than $d$.
Proposition 3 (At most $d$ steps).
For any dictionary in any arrangement of $\mathbb{R}^{d}$, consecutively applying the Zero pivot always attain its terminal in at most $d$ steps.
Proof: According to Lemma 1, it can be concluded that each time the Zero rule is applied, the column it selects will be at least one position later compared to the previous step. In any dictionary in $\mathbb{R}^{d}$, there are a total of $d$ columns, excluding the column corresponding to $x_{g}$. This implies that the Zero rule will be repeated for at most of $d$ times.
@@
The property given in Proposition 3 is very important because most pivot rules, such as Bland’s rule, the Criss-Cross rule, the Random-Edge pivot rule, the Random-Facet pivot rule, converge in exponential or subexponential steps. The exponential convergence of Bland’s rule is shown in [32, 40], while the Criss-Cross rule is shown by an example that requires at least $12^{n}-$ 1 steps for convergence [36]. Moreover, even better method, such as the Random-Edge pivot rule, has an exponential upper bound on pivot steps [18], while the Random-Facet pivot rule in [21] has a subexponential upper bound. In comparison, our Zero rule exhibits a very fast convergence speed. Additionally, for different hyperplane arrangements with the same $d$ and $n$, their upper bounds are different, making their pivot difficult to study. However, our Zero pivot has a common upper bound for each hyperplane arrangement in the same $d$, which is more convenient.
After studying the number of Zero pivots, we focus on the number of terminals in the arrangement. Will there exist multiple terminal dictionaries? To address this question, it is necessary to identify what "0” means in each dictionary.
Lemma 2.
@@
for all $i\in B$ and $j\in N_{\neq g}$. Substituting Eq. (3) into $x_{i}=\sum\limits_{j\in N}\bar{a}_{ij}x_{j}$, we can derive the relationship between $\mathbf{c}_{i}$ and $\mathbf{c}_{j}$ as follows:
@@
From the above lemma, it can be seen that the proposed Zero rule is not an indexing game. Instead, it is deeply related to the properties of the arrangement itself, as represented by the dictionary. The relationship between zero and non-zero entries is the reason why we call our proposed rule the Zero rule. Now, we show that under the Zero rule, there only exists one terminal.
Proposition 4 (Unique terminal).
The collection of terminal dictionaries of any hyperplane arrangement is single.
Proof: Let $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$ be the dictionary such that $N$ has attained the smallest lexicographical order among all co-basis in the hyperplane arrangement; note that the dictionary must be a terminal of the Zero rule. Furthermore, suppose there exists another terminal dictionary, namely $\mathbf{x}_{\tilde{B}}=\tilde{A}\mathbf{x}_{\tilde{N}}$ and $\tilde{N}\neq N$, that is,
@@
Since $N$ is a co-basis in the arrangement, then $\{\mathbf{c}_{l}~{}|~{}l\in N_{0}\}\cup\{\mathbf{c}_{n_{j}}\}\subseteq\{\mathbf{c}_{l}\ ~{}|~{}\ l\in N\}$ is linearly independent, which contradicts with Eq. (7). Hence, in the first place, any dictionary with a co-basis not attaining the lexicographical minimum order is not the end of pivoting by the Zero rule. It implies that the collection of terminal dictionaries of any hyperplane arrangement is single, which is the co-basis with the lexicographical minimum order in the arrangement.
@@
Remark 2: It can be seen that the proposed Zero rule is independent of the column $x_{g}$ which corresponds to the constant terms in the dictionary. This means that when the constant terms in the dictionary change, the selection made by the Zero rule remains unchanged. According to our definition of the dictionary, these constant terms originate from the right-hand side constants of the hyperplanes, namely $b_{1},\cdots,b_{n}$. Changes in these constants geometrically represent the translation of the hyperplanes, but in the dictionary representation, they only affect the column $x_{g}$. Therefore, we can conclude that the Zero pivot is translation invariant.
Connection to an open question. Propositions 3 and 4 might be related to an open question: "Is it possible to find a polynomial pivot rule for linear programming, i.e., a pivot rule which bounds the number of pivot steps by a polynomial function of the number of variables and the number of inequalities, or to prove that such a pivot rule does not exist". This question is important [40], motivated by analyzing the complexity of the simplex method in linear programming. The diameter of a polytope provides a lower bound on the number of steps needed by the simplex method, although it is actually weaker than the so-called Hirsch conjecture in polyhedral theory. The Hirsch conjecture, formulated by W. M. Hirsch in 1957 in [24] and reported in the 1963 book of Dantzig [12], states that the edge-vertex graph of an $n$-facet polytope in $d$-dimensional Euclidean space has a diameter no more than $n-d$. This means that any two vertices of a polytope must be connected by a path of length at most $n-d$. Although there are many studies on the conjecture and diameter of a polytope [6, 28, 35], the conjecture is shown to be true for (0,1)-polytopes [31] and generally false by a counterexample presented in [37] which introduces a 43-dimensional polytope of 86 facets with a diameter exceeding 43.
However, this counterexample does not affect the analysis of the simplex method. It includes the construction of a pivot rule starting from outside the feasible region with an upper bound of $n-d$. The recently developed Facet pivot rule exemplifies the latter possibility. It can achieve an optimal dictionary within at most $n-d$ pivot steps and enters the feasible region from the outside in the final step [45, 46]. Nevertheless, its crucial aspect lies in ensuring dual feasibility at each step, making it difficult to aid the construction of a pivot rule that allows movement between adjacent vertices within the feasible region [45, 46]. In contrast, the Zero rule does not impose stringent conditions in each pivot step. Therefore, our rule, bounded by a linear function of the dimension, may shed some light on the open question. The Zero rule is based on the dictionary itself and goes without a ratio test. Thus, it does not inherently traverse between adjacent vertices or reflect the number of edges between vertices. However, if a ratio test could be incorporated or primal feasibility could be maintained within the Zero rule, it might serve as a polynomial pivoting rule on polytopes.
@@ 4 VE Algorithm Using the Zero Rule
In the above, we introduce the Zero rule. Here, we describe the algorithm proposed by Avis and Fukuda [2], introduce our algorithm using the Zero rule, and conduct a detailed complexity analysis.
4.1 Description of the AF algorithm [2]
The AF algorithm [2] is based on the following principle: within the entire hyperplane arrangement, starting from any given dictionary, the Criss-Cross pivot uniquely provides a finite-length path to an optimal dictionary. All such paths collectively form a graph rooted in this optimal dictionary. Therefore, we can begin from the optimal dictionary and apply valid reverse Criss-Cross pivots in lexicographic order, which corresponds to a depth-first search over the graph. Each dictionary is counted once. By searching every optimal dictionary, we can obtain all dictionaries within the hyperplane arrangement. Meanwhile, to address the situation where a vertex corresponds to multiple dictionaries, we only need to conduct a test when the algorithm discovers a new dictionary to ensure that among all dictionaries associated with the same vertex, only one meets the test criteria, which will then be output or recorded.
Thus, the algorithm can be divided into three steps: first, construct an initial optimal dictionary from the given vertex; perform a reverse search related to the Criss-Cross rule for each optimal dictionary, outputting all dictionaries that meet the testing criteria; after the reverse search, generate another optimal dictionaries based on the currect one and then go second step. If there is no more optimal dictionaries, the algorithm terminate. The key aspects of this algorithm are the testing criteria and the method used to conduct the reverse search. As Figure 4 shows, the testing criterion established by Avis and Fukuda involves selecting the dictionary whose basis is lexicographically minimum among all those corresponding to the same basic solution. For a non-degenerate dictionary associated with a non-degenerate vertex, this dictionary is the only one associated with that vertex; therefore, no need to compare. Regarding the degenerate dictionary associated with a degenerate vertex, multiple degenerate dictionaries corresponding to the same vertex may exist; thus, we need to apply the following standard for selection.
Proposition 5.
Let $B$ be a basis for a degenerate dictionary $\mathbf{x}_{B}=\bar{A}\mathbf{x}_{N}$. $B$ is not lexicographically minimum for the corresponding basic solution if and only if there exists $r\in B_{\neq f}$ and $s\in N_{\neq g}$ such that $r>s,\bar{a}_{rg}=0$ and $\bar{a}_{rs}\neq 0$
Figure 4: An example of the lex-min test. LHS: fails the test. RHS: passes the test.
As for the depth-first search mentioned earlier, since our goal is to find the paths generated by the Criss-Cross pivot, it is necessary to determine whether a pivot $(s,r)$, where $x_{s}\in B$ and $x_{r}\in N$, on a dictionary is a valid reverse Criss-Cross pivot. To do this, the pivot is performed, but we do not use it to update the dictionary that the currently investigated, and the Criss-Cross rule is applied to the resulting dictionary. If the selection yields $(r,s)$, then the pivot $(s,r)$ on the currently investigated dictionary is a valid reverse Criss-Cross pivot. We update the currently investigated dictionary to the computed dictionary and initialize the checking position. Otherwise, do not update and move directly to the next position to repeat the checking process. Additionally, Avis and Fukuda [2] provide a necessary condition to test before performing the pivot, which helps reduce some unnecessary computations during the process. The LHS of Figure 5 illustrates the whole process.
Figure 5: The process of checking and performing valid reverse pivots. LHS: Avis and Fukuda [2]; RHS: Our proposed algorithm.
If every position in the current dictionary has been checked, they then apply the Criss-Cross rule for selection and perform a Criss-Cross pivot, followed by updating the dictionary and the search position. If the Criss-Cross rule does not select a proper position, the reverse search concerning one optimal dictionary is finished, thereby proceeding with another optimal dictionary.
Remark 3. It can be observed that determining whether a pivot is a valid reverse Criss-Cross pivot is quite complex, and sometimes results in $\mathcal{O}(d(n-d)+n)=\mathcal{O}(nd)$ redundant operations. To address this issue, we can construct an if-and-only-if condition to enhance the AF algorithm [2]. Because this is beyond the scope of the proposed Zero rule and our algorithm, we put this if-and-only-if condition and its detailed analysis and comparison in the Appendix, which may be of independent interest to some readers. We denote the algorithm using this if-and-only-if condition as the Enhanced AF. Although its theoretical operational complexity and space complexity are the same as the original AF, namely $\mathcal{O}(n^{2}dv)$ and $\mathcal{O}(nd)$, in practice, the redundant operations in their algorithm introduce several lower-order terms, which can only be neglected given a large $d$ and a larger $n$. Furthermore, even if they can be ignored in theoretical analysis, they still have a considerable impact during actual execution, whereas the Enhanced AF does not have these terms. In all, we think that the Enhanced AF is more efficient.
4.2 Vertex Enumeration with the Zero Rule and Advantage Analysis
Now, we integrate the Zero rule into the algorithmic framework of Avis and Fukuda [2] for the vertex enumeration. First, we note that the core of the AF algorithm is to have a rule to generate a path that leads to an optimal dictionary of the entire hyperplane arrangement. This means that any other pivoting rule including the Zero rule that generates a path towards an optimal dictionary of the entire hyperplane arrangement could also seamlessly fit this framework. Algorithm 1 shows the proposed new algorithm with the Zero rule.
@@ 4.2 Vertex Enumeration with the Zero Rule and Advantage Analysis
Advantage analysis and detailed formulation. Since our focus is on the VE of an arrangement, we only compare with the Criss-Cross rule and Jensen’s general relaxed recursion that are applicable to this problem. The advantages of the Zero rule are as follows:
@@
Jensen’s general relaxed recursion requires partitioning a dictionary and analyzing it case by case. Its termination dictionary could be primal feasible, dual feasible, or optimal [22, 40], which makes it tedious for VE. As for the Criss-Cross rule, although it also needs to discuss two cases, it has only one type of termination dictionary, namely the optimal dictionary. This makes it better than Jensen’s general relaxed recursion. The Criss-Cross rule is the exact one used in the AF algorithm [2].
@@
However, compared to the Zero rule, the Criss-Cross rule is not ideal in both initialization and the reverse search. During the initialization process, the Zero rule is more convenient in generating all dictionaries where the pivoting rule cannot select a position. By Proposition 4, through our way of initialization, the resultant dictionary with co-basis $\{g,1,2,\cdots,d\}$ is always the unique terminal dictionary. Therefore, after applying the Zero rule to the algorithm, the overall enumeration process can be simplified. More favorably, since the terminal dictionary is the dictionary with the smallest lexicographic order co-basis among all dictionaries in the hyperplane arrangement, it can be obtained simply by renaming the hyperplanes. Let the first dictionary we construct satisfy $N=\{g,1,2,\ldots,d\}$ $B=\{d+1,\ldots,n\}$; this automatically be the unique terminal dictionary. While using the Criss-Cross rules requires setting an objective function and generating all optimal dictionaries, which adds an extra step to the initialization process.
When it comes to the reverse search, the Zero rule is also more favorable. According to Proposition 3, the number of pivot steps in the Zero rule is bounded by $d$, whereas that in the Criss-Cross rule is exponential [36, 40]. We introduce the concept of layers into the algorithm. We assign $\mathtt{layer}=0$ to the terminal dictionary, and when a valid reverse Zero pivot is performed, we update $\mathtt{layer}=\mathtt{layer}+1$; conversely, when a Zero pivot is performed, we update $\mathtt{layer}=\mathtt{layer}-1$. Since the number of pivot steps in the Zero rule is bounded by $d$, there are no dictionaries at the $(d+1)$-th layer, meaning that there is no need to do the valid reverse Zero pivots at the $d$-th layer, which can save the cost of checking the valid reverse.
Fortunately, we can adopt the idea from Avis and Fukuda [2] of "computing but not updating”. For dictionaries at the $(d-1)$-th layer, if the pivot $(s,r)$ is a valid reverse Zero pivot, we compute the pivot $(s,r)$ but do not update the dictionary being examined. After determining whether to output the resultant dictionary, we move directly to the next position, thereby bypassing all redundant computations on dictionaries at the $d$-th layer. However, the Criss-Cross rule requires examining every dictionary, which incurs more computational costs.
A graphic explanation of our algorithm is on the RHS of Figure 6, and the pseudocode for it is presented in Algorithm 1.
Figure 6: A graphic explanation of the Enhanced AF (LHS) and ours (RHS).
Algorithm 1 Ours, doing the reverse search using the Zero rule
function $\mathtt{Search}(B,N,\bar{A})$
     $i=1,j=2,\mathtt{layer}=0$;Use $i,j$ represent the $i$-th row, $j$-th column in $\bar{A}$.
     if $\mathtt{lex-min}(B,N,\bar{A})$==1 then
         Print $B$;
     end if
     while $j\leq\mathtt{length}(N)+1$ do
         if $j\leq\mathtt{length}(N)$ then
              if $\mathtt{reverse}(B,N,\bar{A},i,j)==1$ and $\mathtt{layer}<d-1$ then
                  $[B,N,\bar{A}]=\mathtt{pivot}(B,N,\bar{A},i,j)$; % Compute and update.
                  if $\mathtt{lex-min}(B,N,\bar{A})==1$ then
                       Print $B$;
                  end if
                  $i=1,j=2,\mathtt{layer}=\mathtt{layer}+1;$Initialize the position, go next layer.
              elseif   $\mathtt{reverse}(B,N,\bar{A},i,j)==1$ and $\mathtt{layer}==d-1$
                  $[tempB,tempN,temp\bar{A}]=\mathtt{pivot}(B,N,\bar{A},i,j)$; % Not update.
                  if $\mathtt{lex-min}(tempB,tempN,temp\bar{A})==1$ then
                       Print $tempB$;
                  end if
                  $[i,j]=\mathtt{increment}(i,j)$;Go next position.
              else
                  $[i,j]=\mathtt{increment}(i,j)$;Invalid, go next position.
              end if
         elseEach position on this dictionary has been checked.
              $[i,j]=\mathtt{select}(B,N,\bar{A})$;Use Zero rule to select a position.
              if both $i$ and $j$ nonempty thenThe Zero rule find a position.
                  $[B,N,\bar{A}]=\mathtt{pivot}(B,N,\bar{A},i,j)$;Back to parent dictionary.
                  $[i,j]=\mathtt{find}(i,j)$;Find the corresponding position.
                  $[i,j]=\mathtt{increment}(i,j)$;
                  $\mathtt{layer}=\mathtt{layer}-1$;
              elseThe Zero rule does not find a position.
                  Break;The function $Search$ terminated.
              end if
         end if
     end while
end function function $\mathtt{lex-min}(B,N,\bar{A})$Test if the current basis output.
function $\mathtt{reverse}(B,N,\bar{A},i,j)$Test if pivot$(B(i),N(j))$ a valid reverse pivot.
function $[B,N,\bar{A}]=\mathtt{pivot}(B,N,\bar{A},i,j)$Compute the new dictionary.
function $[i,j]=\mathtt{increment}(i,j)$Go next position on the dictionary.
function $[i,j]=\mathtt{select}(i,j)$Use the Zero rule to find a position.
function $[i,j]=\mathtt{find}(i,j)$Find corresponding position after Zero pivot.